### v5

@@ 6.2. Variations on a Theme of Queues
Next, we repair functions and proofs across a change in implementation of a queue data structure. This is motivated by an example from Angiuli et al. (2021), which showed that quotient types can be use to adjust certain relations more general than equivalences into equivalences for use with transport in Cubical Agda. That class of changes was cited in the Pumpkin Pi paper as an example that could not be expressed naturally in Coq with the original framework. With our extensions to Pumpkin Pi, we can express this using setoids. Our Coq implementation of this case study can be found in two_list_queue_equivalence_repair_tool.v.
@@
We now use our extension to repair functions across this change. We provide the standard queue API by repairing enqueueOLQ and dequeueOLQ, as well as the helper functions dequeueHelpOLQ and returnOrEnqTLQ.
@@
We prove a theorem dequeueEnqueueOLQ stating that enqueue and dequeue commute in the expected way, using returnOrEnq as a helper function. We repair this automatically, with none of the workarounds from the previous case study, since the proof of dequeueEnqueueOLQ is defined using the option_rect rather than depElimProp. We also repair the proof of dequeueEmptyOLQ, providing an algebraic specification for the repaired datatype.
@@
forall (exp : nat), coeff p1 exp = coeff p2 exp.
Figure 11. Definitions and equivalence relations for CLPoly and CEPPoly. removeLeadingZeros l removes leading zeros from l, while coeff p exp is the nth degree coefficient of p.
Definition depConstr (l : opaque_list) (p : noLeadingZeros l) : CLPoly.
Definition depRec (C : Type) (X : forall (l : opaque_list) (p : noLeadingZeros l), C) (p : CLPoly) : C.
Figure 12. The types for depConstr and depRec for CLPoly, where opaque_list is an alias for list nat.
@@
As in the previous case study, we have finished repairing proofs, but our repaired implementation of dequeue is inefficient. Thus, we implement the fast version of dequeue descirbed earlier, and follow the same methodology as in the previous case study to port our repaired proofs to use this fast version. We first show pointwise equality of the functions, and then rewrite inside of our proofs, giving us repaired proofs about the fast dequeue function.
@@ 6.3. Polynomial Polynomials
For our third case study, we provide sparse and dense representations of univariate polynomials with natural number coefficients, both of which are represented using setoids. We implement and repair addition of polynomials and evaluation of polynomials on a natural number, as well as proofs that addition is commutative and that evaluation respects addition. Our proofs can be found in polynomial.v.
@@
Our first representation of polynomials, CLPoly (short for coefficient list polynomial), is as lists of natural numbers. The members of the list are the coefficients of the polynomial in order of decreasing degree. Two CLPolys are equivalent if they are equal after removing leading zeros. Our second representation, CEPPoly (short for coefficient-exponent pair polynomial), is lists of pairs of natural numbers. Each pair (c, exp) represents a monomial in a sum, with c the coefficient and exp the exponent of the monomial, with duplicate exponents allowed. Two members of CEPPoly are equivalent if the polynomial they represent has the same coefficients. Both representations are in Figure 11.
@@
The types of depConstr for both CLPoly and CEPPoly take the shape of the constructor of this type, as do the types of the eliminators and $\iota$-reduction rules. We show this for part of our configuration for CLPoly in Figure 12.
@@
We use CLPoly as our source setoid and CEPPoly as our target setoid. We define add and eval, representing addition and evaluation, explicitly annotated with the components of the configuration. We automatically repair these functions using our extension to Pumpkin Pi. Our extension automatically proves that eval is proper, but fails to show add is proper, so we do so manually.
@@
We repair proofs of addComm, which states that add is commutative, and evalRespectsAdd, which states eval (add p1 p2) n = (eval p1 n) + (eval p2 n). Both use depElimProp, so we reconfigure Pumpkin Pi to use specialized versions of depElimProp, as we did in the first case study. From there, both proofs repair automatically. Neither of these proofs use rewriting. Thus, to demonstrate lifting setoid rewrites to setoid rewrites, we also repair three simple proofs about polynomials specifically using setoid rewriting.
@@
The constructor for CLPoly, seen in Figure 12, takes an argument of type opaque_list. Because CLPoly is internally list nat, Pumpkin Pi will attempt to repair all instances of list nat. However, we define functions over list nat which we do not want to repair. Thus, we define an alias opaque_list for list nat and tell Pumpkin Pi not to repair opaque_list or any functions and theorems that should not be repaired. This behavior and our strategy for dealing with it are inherited from Pumpkin Pi.
7. Discussion: Grounding in Cubical
We extended Pumpkin Pi with additional automation for setoids in order to support a new class of changes corresponding to quotient type equivalences. This is in spirit with Pumpkin Pi’s existing philosophy, as Pumpkin Pi is a tool that is external to Coq. But, to make sense of Pumpkin Pi’s behavior, we need a univalent metatheory. This applies for both the existing behavior and for our extensions.
This discussion goes further than prior work on Pumpkin Pi, grounding the meaning not just in an uninstantiated univalent metatheory, but also directly in Cubical Agda. In particular, we demonstrate how setoids approximate quotient types (Section 7.1) as well the tradeoffs between those representations. For the original Pumpkin Pi work, this lets us formalize, for the first time ever, some of the correctness criteria for repair (Section 7.2). Finally, it reveals that the ad hoc treatment of replacing slow implementations with fast implementations in Coq could become much more grounded as well (Section 7.3). To concretize this discussion, we manually reimplement two case studies from Section 6 in Cubical Agda, which we use throughout this section as examples.
7.1. Grounding Setoids
Because Coq lacks quotient types, our implementation uses setoids to represent quotients. This incurred additional implementation overhead. To understand this overhead and where it manifests, we revisit the same proof repair problem manually in Cubical Agda, which has native quotient types.
In Cubical Agda, set quotient types are encoded internally as higher inductive types:
@@ 7.1. Grounding Setoids
For example, consider the equivalence between the two integer types from Figure 6 in Section 6.1. In Cubical Agda, the type GZ is defined as ($\mathbb{N}$$\times$ $\mathbb{N}$)$\sim$, where $\sim$ represents our equivalence relation. The resultant type is the quotient of $\mathbb{N}$ by $\sim$. Since quotient types are internal, a quotient type equivalence in Cubical Agda is just a normal type equivalence. See grothendieck_int_equivalence.agda.
In Agda, we must prove that our depElims are proper functions upon definition; then, all functions defined using them are automatically proper. We may skip this proof when defining depRec when using setoids in Coq, but the burden appears when we repair proofs. When using setoids, we can only rewrite under functions we have proven proper. By proving that depRec is proper, our automation handles many of these proof obligations, but not every case as in Agda.
@@
More consequentially, the additional rules to repair equivalence relations from Figure 4 in Section 4 are not needed at all in a restricted set fragment of Cubical Agda. The transformation from Figure 3 in Section 4 works, assuming the same annotations as in Coq, noting that we cannot reuse any of Pumpkin Pi’s automation in Cubical Agda). There is only one additional restriction: every motive P : GZ $\to$ Set comes with the requirement that ((x : GZ) $\to$ isSet (P x)), to ensure that we actually do stay in the set fragment.
To demonstrate this, we manually followed the transformation to repair the functions and proofs from the first two case studies, which can be found in the supplementary materials in files grothendieck_int_equivalence.agda and
two_list_queue_equivalence.agda respectively.
7.2. Grounding Correctness
We would like to prove that our functions and proofs were repaired correctly. As noted in previous work (Ringer et al., 2021; Ringer, 2021), stating this relies on a univalent notion of equality: we want to know that terms behave the same up to the equivalence we repair across. Without assuming univalence as an axiom in Coq (which goes against the very philosophy of Pumpkin Pi), we cannot state these correctness criteria internally.
This was done metatheoretically in a univalent type theory for the Pumpkin Pi transformation that we build off of, but it was not proven for any particular example at all. In Cubical Agda, heterogeneous path equalities in the form of PathP types give us the power needed to formalize the correctness theorems about repair from the Pumpkin Pi paper in Cubical Agda, and even to prove them correct on an example type equivalence for the first time.
@@ 7.2. Grounding Correctness
Some of these theorems are generic across all types. For example, we internally prove correctness of the Lam rule of the transformation from Figure 3. Other rules are stated specifying the types being repaired. For example, we proved the repaired eliminator we defined for a simple quotient equivalence was correct. Our source type was $\mathbb{N}$, and our target was Int / rInt, where Int = $\mathbb{N}\uplus\mathbb{N}$ and rInt relates inl n and inr n. The dependent constructors and eliminators correspond to the usual ones for $\mathbb{N}$. The correctness condition for a repaired eliminator for a given configuration was stated externally in the Pumpkin Pi paper, but it was not proven for any type. We adapted this theorem to Cubical Agda for our example and, for the first time, proved that it held. For details, see equivalence_int_abs.agda in the supplementary material, or Figure 13 and Figure 14 in the appendix.
Even when we can internally prove individual rules of the transformation correct, it can be challenging to actually compose those proofs to get proofs of the correctness of particular repaired proofs. Our correctness theorem is stated in terms of PathPs, but showing that two PathPs are equal requires reasoning about the equality of equality proofs between types. This means entering the fragment of cubical that is proof relevant: identity proofs need not be unique, and specific proofs may be needed for goals. Concretely, we were able to compose the correctness proofs to show the correctness of repaired functions, like addition. But we could not always do the same for many more interesting proofs, like commutativity of addition, even after proving full correctness of the eliminator. This is why we say that we are able to show just some correctness proofs internally in Cubical Agda, even though in theory, it should be possible to show all of them. Fully reckoning with proof relevance so that correctness proofs compose correctly requires additional work. For this reason, we have yet to successfully implement internal correctness proofs in the Cubical Agda versions of the case studies we discussed in Section 6.
7.3. Grounding Fast and Slow
Recall that two of the case studies in Coq used an ad hoc process for moving between fast and slow implementations of repaired functions. This is due to the lack of functional extensionality and internal transport in Coq. In Cubical Agda, functional extensionality and transport are both theorems. This makes the process of moving between slow and fast implementations of repaired functions straightforward, grounding it in computational notions of functional extensionality and transport. It also shows that even in a world with external proof repair, internal transport can still be quite useful.
Consider the addition example from Section 6.1. In Cubical Agda, we call our fast addition function addGZ, and our slow repaired addition function _+GZ_. Using functional extensionality, we can show that these two functions are path-equal to each other. Then, to obtain proofs for our fast addition function, we substitute the fast function for the repaired function in the repaired proofs using transport. The proofs for this can be found in grothendieck_int_equivalence.agda in the supplementary material, and similar proofs for the two list queue example are in two_list_queue_equivalence.agda.

### v6

@@ 5.2. Variations on a Theme of Queues
Next, we repair functions and proofs across a change in implementation of a queue data structure. This is motivated by an example from Angiuli et al. (2021), which showed that quotient types can be used to adjust certain relations more general than equivalences into equivalences for use with transport in Cubical Agda. That class of changes was cited in the Pumpkin Pi paper as an example that could not be expressed naturally in Coq with the original framework. With our extensions to Pumpkin Pi, we can express this using setoids. Our Coq implementation of this case study can be found in two_list_queue_equivalence_repair_tool.v.
@@
We now use our extension to repair functions across this change. We provide the standard queue API by repairing enqueueOLQ and dequeueOLQ, as well as the helper functions dequeueHelpOLQ and returnOrEnqTLQ. The first three can be found in Figure 13, and the last in Figure 14.
@@
We prove a theorem dequeueEnqueueOLQ, found in Figure 14, stating that enqueue and dequeue commute in the expected way, using returnOrEnq as a helper function. We repair this automatically, with none of the workarounds from the previous case study, since the proof of dequeueEnqueueOLQ is defined using the option_rect rather than depElimProp. We also repair the proof of dequeueEmptyOLQ, providing an algebraic specification for the repaired datatype.
Definition fastDequeueTLQ (q : TLQ) :=
let (l1, l2) := q in match l1, l2 with
| [] , [] => None
| h1 :: t1 , [] =>
Some (([] , tl (rev l1)), hd h1 (rev l1))
| _ , h2 :: t2 => Some ((l1, t2), h2)
end.
Figure 15. Fast dequeue function for two list queues.
@@
forall (exp : nat),
coeff p1 exp = coeff p2 exp.
Figure 16. Definitions and equivalence relations for CLPoly and CEPPoly. removeLeadingZeros l removes leading zeros from l, while coeff p exp is the nth degree coefficient of p. We provide instances of Equivalence eq_CLPoly and Equivalence eq_CEPPoly in the supplementary material.
Definition depConstrCLPoly
(l : opaque_list)
(p : noLeadingZeros l) : CLPoly.
Definition depRecCLPoly (C : Type)
(X : forall (l : opaque_list)
(p : noLeadingZeros l), C)
(p : CLPoly) : C.
Definition depConstrCEPPoly
(l : list nat)
(p : noLeadingZeros l) : CEPPoly.
Definition depRecCEPPoly (C : Type)
(X : forall (l : list nat)
(p : noLeadingZeros l), C)
(p : CEPPoly) : C.
Figure 17. The types of depConstr and depRec for CLPoly on the left and CEPPoly on the right, where opaque_list is an alias for list nat.
Definition addCLPoly (p1 p2 : CLPoly) :=
depRecCLPoly CLPoly
(fun (l : opaque_list)
(p : noLeadingZeros l) =>
depRecCLPoly CLPoly
(fun (l0 : opaque_list)
(p0 : noLeadingZeros l0) =>
depConstrCLPoly (addLists l l0)
(addListsNoLeadingZeros l
l0 p p0))
p2)
p1.
Definition evalCLPoly (p : CLPoly)
(n : nat) :=
depRecCLPoly nat
(fun (l : opaque_list)
(proof : noLeadingZeros l) =>
evalList l n)
p.
Definition addCEPPoly (p1 p2 : CEPPoly) :=
depRecCEPPoly CEPPoly
(fun (l : opaque_list)
(p : noLeadingZeros l) =>
depRecCEPPoly CEPPoly
(fun (l0 : opaque_list)
(p0 : noLeadingZeros l0) =>
depConstrCEPPoly (addLists l l0)
(addListsNoLeadingZeros l
l0 p p0))
p2)
p1.
Definition evalCEPPoly (p : CEPPoly)
(n : nat) :=
depRecCEPPoly nat
(fun (l : opaque_list)
(proof : noLeadingZeros l) =>
evalList l n)
p.
Figure 18. The definitions of addition and evaluation for polynomials. The original, over CLPoly, is on the left, and the repaired version over CEPPoly is on the right.
Theorem addComm :
forall (p1 p2 : CLPoly),
eq_CLPoly (add p1 p2) (add p2 p1).
Theorem evalRespectsAdd :
forall (p1 p2 : CLPoly) (n : nat),
eval (add p1 p2) n =
(eval p1 n) + (eval p2 n).
Figure 19. The types of addComm and evalRespectsAdd for CLPoly. These theorems were repaired to CEPPoly.
@@
As in the previous case study, we have finished repairing proofs, but our repaired implementation of dequeue is inefficient. Thus, we implement the fast version of dequeue described earlier, found in Figure 15, and can follow the same methodology as in the previous case study to port repaired proofs to use this fast version. We first show pointwise equality of the functions, and then rewrite inside of our proofs, giving us repaired proofs about the fast dequeue function.
@@ 5.3. Polynomial Polynomials
For our third case study, we provide sparse and dense representations of univariate polynomials with natural number coefficients, both of which are represented using setoids. The former representation uses the simplest possible data type to represent a polynomial, while the latter representation will mimic the way polynomials are commonly written, which will utilize far less storage for polynomials of high degree with many terms having a coefficient of 0. We implement and repair addition of polynomials and evaluation of polynomials on a natural number, as well as proofs that addition is commutative and that evaluation respects addition. Our proofs can be found in polynomial.v.
@@
Our first representation of polynomials, CLPoly (short for coefficient list polynomial), is as lists of natural numbers. The members of the list are the coefficients of the polynomial in order of decreasing degree. For example, $x^{2}+3$ is represented as [1; 0; 3]. Two CLPolys are equivalent if they are equal after removing leading zeros. Our second representation, CEPPoly (short for coefficient-exponent pair polynomial), is lists of pairs of natural numbers. Each pair (c, exp) represents a monomial in a sum, with c the coefficient and exp the exponent of the monomial, with duplicate exponents allowed. Using the same example, $x^{2}+3$ would be represented as [(1, 2); (3, 0)]. Two members of CEPPoly are equivalent if the polynomial they represent has the same coefficients. Both representations are found in Figure 16.
@@
The types of depConstr for both CLPoly and CEPPoly take the shape of the constructor of this type, as do the types of the eliminators and $\iota$-reduction rules. We show this for part of our configuration for CLPoly and CEPPoly in Figure 17.
@@
We use CLPoly as our source setoid and CEPPoly as our target setoid. We define add and eval, representing addition and evaluation, explicitly annotated with the components of the configuration. We automatically repair these functions using our extension to Pumpkin Pi. The original definitions and the repaired versions can be found in Figure 18. Our extension automatically proves that eval is proper, but fails to show add is proper, so we do so manually.
@@
We repair proofs of addComm, which states that add is commutative, and evalRespectsAdd, which states that evaluation distributes over addition. The statements of these theorems can be found in Figure 19. Both use depElimProp, so we reconfigure Pumpkin Pi to use specialized versions of depElimProp, as we did in the first case study. From there, both proofs repair automatically. Neither of these proofs use rewriting. Thus, to demonstrate lifting setoid rewrites to setoid rewrites, we also repair three simple proofs about polynomials which specifically use setoid rewriting.
@@
The constructor for CLPoly, seen in Figure 17, accepts an argument of type
opaque_list. Because CLPoly is internally list nat, Pumpkin Pi will attempt to repair all instances of list nat. However, we define functions over list nat which we do not want to repair. Thus, we define an alias opaque_list for list nat and tell Pumpkin Pi not to repair opaque_list, as well as any other functions and theorems that should not be repaired. This behavior and our strategy for dealing with it are inherited from Pumpkin Pi.
6. Correctness
Up to now, we have been imprecise with what it means for repair to be conducted correctly. The original Pumpkin Pi paper (Ringer et al., 2021) outlines a definition for correctness of repair using univalent type theory. We review that definition here, giving the needed background in univalent type theory. Then, we implement this definition in Cubical Agda, a proof assistant which has the necessary univalent type theory. This allows us to, for the first time, construct proofs that functions and theorems were correctly repaired.
To implement correctness of repair, we need a notion of heterogeneous equality. The Pumpkin Pi paper uses dependent path equality in a univalent type system for this purpose. Univalence states that equivalence is equivalent to equality (Cohen et al., 2018; Angiuli et al., 2017): any equivalence of types corresponds to a unique proof of equality between those two types. We term these equalities path equalities. Given a path equality p : A $\equiv$ B between two types and given elements a : A, b : B of those types, we can construct the type of dependent path equalities between those elements, notated PathP p a b. This type is inhabited if and only if the equivalence of types corresponding to p by univalence maps a to b.
The repair transformation is parametrized by an equivalence f : A $\to$ B between the old and new types. By univalence, this equivalence gives an equality between those types. This equality of types in turn extends to equalities of types built using those types: for instance, if A $\equiv$ B, then A $\to$ C $\equiv$ B $\to$ C. For any repaired term, then, there is an equality between the type of the original term and the type of the repaired term. We obtain this equality inductively in much the same way we do repair inductively. Whenever an inductive type appears in the old term, if that type is A, we use the equality proof corresponding to f. Whenever another inductive type appears, we use refl as the equality proof. When the other rules would apply, we have theorems constructing an equality proof from the proofs for the constituent terms. Thus, we can construct the type of dependent path equalities between these terms. If that type is inhabited, we say that the term was repaired correctly.
Thus, to use this approach to construct proofs of correct repair, we need to work in a type system that supports this notion of dependent path equality. Coq does not support univalence without adding axioms to its type system. Instead, we implement this approach using Cubical Agda, which has univalence as a theorem. As a technical detail, we restrict ourselves to working with types which are h-sets: that is, types where uniqueness of identity proofs holds. As a bonus, Cubical Agda allows a direct implementation of quotient types via the following higher inductive type:
@@ 6. Correctness
The first constructor is the constructor for an element of a quotient type as described in Section 2, and the second encodes the equality property for quotient types. The third constructor enforces that quotient types are h-sets. We will use this type instead of setoids when specifying correctness of repair in Cubical Agda. As a result, the additional rules to repair equivalence relations from Figure 5 in Section 3 are not needed when conducting repair in Cubical Agda. The transformation from Figure 4 in Section 3 works, assuming the same annotations as in Coq, noting that we cannot reuse any of Pumpkin Pi’s automation in Cubical Agda. There is only one additional restriction: for a quotient type Q, every motive P : Q $\to$ Set comes with the requirement that ((x : Q) $\to$ isSet (P x)), to ensure that we actually do stay in the h-set fragment of Cubical Agda. To demonstrate that our repair methodology still works under this paradigm, we manually followed the transformation to repair the functions and proofs from the first two case studies in Section 5, which can be found in the supplementary materials in files grothendieck_int_equiv.agda and equivalence_queue.agda respectively.
lamOK: {T} {F}
(f: (t: T i0) $\to$ F i0 t) (f’: (t: T i1$\to$ F i1 t)
[…]
PathP ($\lambda$ i $\to$ $\forall$ (t : T i) $\to$ F i t) f f
[…]
Figure 20. A theorem showing that the Lam rule is correct. Here, i, i0, and i1 are terms of the interval type, which is a primitive construct in cubical used to define path equalities. The rest is analogous to Figure 4: f is the left function in the rule, f is the right function, F i0 is the type of f, F i1 is the type of f, and all other subterms have the same names.
Then, we go about proving theorems which state that each repair rule repairs terms correctly, given that the inputs to the rule are themselves correctly repaired. Some of these theorems are generic across all types. For example, we internally prove correctness of the Lam rule of the transformation from Figure 4 in Figure 20, which is generic across any repair instance.
elimOK :
[…]
$\forall$ (PA : $\mathbb{N}$to Type) (PB : Int / rInt $\to$ Type) (PBSet : $\forall$ b $\to$ isSet (PB b)) $\to$
[…]
PathP ($\lambda$ i $\to$ Type) (PA a) (PB b)) $\to$
$\forall$ (PAO : PA zero) (PBO : PB depConstrInt/rInt0) $\to$
[…]
$\forall$ (PAS : $\forall$ a $\to$ PA a $\to$ PA (suc a)) (PBS : $\forall$ b $\to$ PB b $\to$ PB (depConstrInt/rIntS b)) $\to$
[…]
(PAS a IHa)
(PBS b IHb)) $\to$
[…]
(Nat.elim {A = PA} PAO PAS a)
(depElimSetInt/rInt PB PBSet PBO PBS b)
Figure 21. The theorem stating the correctness condition for the repaired dependent eliminator for a simple example type, which has been proven internally in Cubical Agda. This theorem shows that, if all the inputs to the eliminator correspond to each other across the isomorphism, then the output of the eliminator applications also corresponds across that isomorphism. Here, depConstr0OK and depConstrSOK are the correctness proofs of the repaired constructors, also proven internally.
@@
Other rules are stated specifying the types being repaired. For example, we proved the repaired eliminator we defined for a simple quotient equivalence was correct. Our source type was $\mathbb{N}$, and our target was Int / rInt, where Int = $\mathbb{N}\uplus\mathbb{N}$ and rInt relates inl n and inr n. The dependent constructors and eliminators correspond to the usual ones for $\mathbb{N}$. The correctness condition for a repaired eliminator for a given configuration was stated externally in the Pumpkin Pi paper, but it was not proven for any type. We adapted this theorem to Cubical Agda for our example and, for the first time, proved that it held. The type of the rule showing that the dependent eliminator repairs correctly can be found in Figure 21.
Using these rules, we were able to compose the correctness proofs to show the correctness of repaired functions, like addition:
addCorrect : $\forall$ (a b : $\mathbb{N}$) (a b : Int / rInt) $\to$
[…]
We also were able to prove theorems about these functions. For instance, below is the type of a term proving that our proof that addition is commutative on $\mathbb{N}$, addCommNat, was correctly repaired to a proof that addition is commutative on our new type:
addCommCorrect :
[…]
PathP ($\lambda$ i $\to$ addCorrect a b a b pa pb i $\equiv$ addCorrect b a b a pb pa i)
(addCommNat a b) (addCommInt/rInt a b’)
When constructing proofs of correct repair for theorems, because of difficulties when composing different PathPs, our rules for proving correct repair do not always compose directly in their current formulation. As a result, if Cubical Agda had the facilities for building automation, we would not currently have a complete procedure for constructing proofs of correct repair fully automatically. However, we are able to prove these theorems directly in such cases, demonstrating that our repair procedure produces the desired output. All of the rules we prove, as well as the proofs of correct repair in our example, can be found in equivalence_int_abs.agda in the supplementary material.